# Complexity Analysis of Normalizing Constant Estimation: from Jarzynski Equality to Annealed Importance Sampling and beyond

### v1

@@ H.1 Thermodynamic Integration
We first briefly review the thermodynamic integration (TI) algorithm. Its essence is to write the free-energy difference as an integral of the derivative of free energy. Consider the general curve of probability measures $(\pi_{\theta})_{\theta\in[0,1]}$ defined in Eq. 11. Then,
@@
One may choose time points $10=\theta_{0}<...<\theta_{M}=$ 1 and approximate Eq. 36 by a Riemann sum:
@@
where the expectation under each $\pi_{\theta_{\ell}}$ can be estimated by sampling from $\pi_{\theta_{\ell}}$. Nevertheless, there is a way of writing the exact equality instead of the approximation in Eq. 37: since
@@
We follow the notations used in Thm. 7 to demonstrate the following claim: TI (Eq. 38) is a special case of AIS with every transition kernel $F_{\ell}(x,\cdot)$ chosen as the perfect proposal $\pi_{\theta_{\ell}}$.
@@
and hence the AIS equality becomes the following identity, exactly the same as Eq. 36:
@@
(III) Distinction between equilibrium and non-equilibrium methods.
@@
H.2 Proof of the Second Part of Lem. 19
Recall that our goal is to estimate $\pi_{0}$’s normalizing constant $Z_{0}=\int\mathrm{e}^{-V_{0}}\mathrm{d}x$, where $V_{0}$ is $\beta$-strongly convex and $3\beta$-smooth, with global minimizer $x^{\prime}$ satisfying $\|x^{\prime}\|\leq R\lesssim\frac{1}{\sqrt{\beta}}$. The aim is to obtain an estimator $\widehat{Z}_{0}\approx Z_{0}$ such that
$\operatorname{Pr}(\mathcal{F})\leq\frac{1}{8},~{}\text{where}~{}\mathcal{F}:=\left\{\left|\frac{\widehat{Z}_{0}}{Z_{0}}-1\right|\geq\frac{\varepsilon}{8}\right\}.$ (40)
Following the discussion above, the TI algorithm goes as follows. Consider a sequence of non-negative numbers $\lambda_{0}>\lambda_{1}>...>\lambda_{K}=0$, where there exists a common $\gamma_{0}>0$ such that $\lambda_{k}=\left(1+\gamma_{0}\right)\lambda_{k+1}$, for all $k\in\left[\!\left[0,K-2\right]\!\right]$. Let $\rho_{k}:=\frac{1}{\zeta_{k}}\mathrm{e}^{-f_{k}}$, where $f_{k}:=V_{0}+\frac{\lambda_{k}}{2}\|\cdot\|^{2}$ is $(\beta+\lambda_{k})$-strongly-convex and $(3\beta+\lambda_{k})$-smooth. One can write
$Z_{0}=\zeta_{K}=\zeta_{0}\prod_{k=0}^{K-1}\underbrace{\frac{\zeta_{k+1}}{\zeta_{k}}}_{=:G_{k}},~{}\text{where}~{}G_{k}=\operatorname{\mathbb{E}}_{\rho_{k}}\underbrace{\exp\left(\frac{\lambda_{k}-\lambda_{k+1}}{2}\|\cdot\|^{2}\right)}_{=:g_{k}},$
and estimate each $G_{k}$ by
$\widehat{G}_{k}:=\frac{1}{N}\sum_{n=1}^{N}g_{k}(\widehat{X}_{n}^{(k)}),~{}\widehat{X}_{n}^{(k)}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\widehat{\rho}_{k}\approx\rho_{k},$
so the final estimator is $\widehat{Z}_{0}:=\widehat{\zeta}_{0}\prod_{k=0}^{K-1}\widehat{G}_{k}$, in which $\widehat{\zeta}_{0}\approx\zeta_{0}$. To proceed, we first prove the following lemma.
Lemma 36.
If
  1. 1.
    $\operatorname{TV}(\widehat{\rho}_{k},\rho_{k})\leq\delta\asymp\frac{1}{NK}$, for all $k\in\left[\!\left[0,K-1\right]\!\right]$.
  2. 2.
    The estimate $\widehat{\zeta}_{0}$ satisfies $\left|\log\frac{\widehat{\zeta}_{0}}{\zeta_{0}}\right|\lesssim\varepsilon$.
  3. 3.
    For all $k\in\left[\!\left[0,K-1\right]\!\right]$, the following equation holds:
    $\frac{\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}^{2}}{\left(\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}\right)^{2}}\leq 1+O(1).$ (41)
Then with $N\asymp\frac{K}{\varepsilon^{2}}$, Eq. 40 holds.
Proof.
By definition of TV distance, for each pair of $(n,k)$ one can construct a random variable $X_{n}^{(k)}\sim\rho_{k}$ that only depends on $\widehat{X}_{n}^{(k)}$ and satisfies $\operatorname{Pr}\left(\widehat{X}_{n}^{(k)}\neq X_{n}^{(k)}\right)\leq\delta$. Define the event
=$\mathcal{E}=\left\{\widehat{X}_{n}^{(k)}=X_{n}^{(k)}:\forall n\in\left[\!\left[1,N\right]\!\right],k\in\left[\!\left[0,K-1\right]\!\right]\right\}.$
By independence, $\operatorname{Pr}(\mathcal{E})\geq(1-\delta)^{NK}\geq 1-\delta NK\gtrsim 1$. If $\operatorname{Pr}(\mathcal{F}|\mathcal{E})\leq\frac{1}{16}$ and $\operatorname{Pr}(\mathcal{E}^{\complement})\leq\frac{1}{16}$, then
$\operatorname{Pr}(\mathcal{F})=\operatorname{Pr}(\mathcal{F}|\mathcal{E})\operatorname{Pr}(\mathcal{E})+\operatorname{Pr}(\mathcal{F}|\mathcal{E}^{\complement})\operatorname{Pr}(\mathcal{E}^{\complement})\leq\operatorname{Pr}(\mathcal{F}|\mathcal{E})+\operatorname{Pr}(\mathcal{E}^{\complement})\leq\frac{1}{8},$
as desired.
To obtain $\operatorname{Pr}(\mathcal{F}|\mathcal{E})\leq\frac{1}{16}$, from now on we always assume that $\mathcal{E}$ happens, and omit the conditional notation $(\cdot|\mathcal{E})$ in probability and expectation for simplicity. Note that in this case, $\widehat{G}_{k}=\frac{1}{N}\sum_{n=1}^{N}g_{k}(X_{n}^{(k)})$, $X_{n}^{(k)}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\rho_{k}$, so $\operatorname{\mathbb{E}}\widehat{G}_{k}=G_{k}$. One can upper bound the probability of large relative error as follows, leveraging Lem. 27:
$\operatorname{Pr}\left(\left|\frac{\widehat{Z}_{0}}{Z_{0}}-1\right|\gtrsim\varepsilon\right)$ $\leq\operatorname{Pr}\left(\left|\log\frac{\widehat{Z}_{0}}{Z_{0}}\right|\gtrsim\varepsilon\right)=\operatorname{Pr}\left(\left|\log\frac{\widehat{\zeta}_{0}}{\zeta_{0}}+\log\prod_{k=0}^{K-1}\frac{\widehat{G}_{k}}{G_{k}}\right|\gtrsim\varepsilon\right)$
$\leq\operatorname{Pr}\left(\left|\log\frac{\widehat{\zeta}_{0}}{\zeta_{0}}+\log\prod_{k=0}^{K-1}\frac{\widehat{G}_{k}}{G_{k}}\right|\gtrsim\varepsilon\right)$
$\leq\operatorname{Pr}\left(\left|\log\prod_{k=0}^{K-1}\frac{\widehat{G}_{k}}{G_{k}}\right|\gtrsim\varepsilon\right)\leq\operatorname{Pr}\left(\left|\prod_{k=0}^{K-1}\frac{\widehat{G}_{k}}{G_{k}}-1\right|\gtrsim\varepsilon\right)$
$\lesssim\frac{1}{\varepsilon^{2}}\operatorname{\mathbb{E}}\left(\prod_{k=0}^{K-1}\frac{\widehat{G}_{k}}{G_{k}}-1\right)^{2}=\frac{1}{\varepsilon^{2}}\left(\prod_{k=0}^{K-1}\frac{\operatorname{\mathbb{E}}\widehat{G}_{k}^{2}}{G_{k}^{2}}-1\right),$
where the last line is due to Markov inequality. Choosing $N\asymp\frac{K}{\varepsilon^{2}}$ yields
$\frac{\operatorname{\mathbb{E}}\widehat{G}_{k}^{2}}{G_{k}^{2}}-1=\frac{\operatorname{Var}\widehat{G}_{k}^{2}}{G_{k}^{2}}=\frac{\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}^{2}-\left(\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}\right)^{2}}{N\left(\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}\right)^{2}}\lesssim\frac{1}{N},$
which implies
$\operatorname{Pr}\left(\left|\frac{\widehat{Z}_{0}}{Z_{0}}-1\right|\gtrsim\varepsilon\right)$ $\lesssim\frac{1}{\varepsilon^{2}}\left(\prod_{k=0}^{K-1}\frac{\operatorname{\mathbb{E}}\widehat{G}_{k}^{2}}{G_{k}^{2}}-1\right)\leq\frac{1}{\varepsilon^{2}}\left(\left(1+\frac{1}{N}\right)^{K}-1\right)\lesssim\frac{K}{N\varepsilon^{2}}\lesssim 1.$
The following lemmas show how to accurately estimate $\zeta_{0}$ and how to satisfy Eq. 41.
Lemma 37.
With $\lambda_{0}\asymp\frac{d\beta}{\varepsilon}$, $\widehat{\zeta}_{0}:=\exp\left(-V_{0}(0)+\frac{\|\nabla V_{0}(0)\|^{2}}{2(3\beta+\lambda_{0})}\right)\left(\frac{2\pi}{3\beta+\lambda_{0}}\right)^{\frac{d}{2}}$ satisfies $\left|\log\frac{\widehat{\zeta}_{0}}{\zeta_{0}}\right|\lesssim\varepsilon$.
Proof.
By assumption, $f_{0}$ is $(\beta+\lambda_{0})$-strongly-convex and $(3\beta+\lambda_{0})$-smooth. Using quadratic upper and lower bounds on $f_{0}$,
$\exp\left(-f_{0}(0)+\frac{\|\nabla f_{0}(0)\|^{2}}{2(3\beta+\lambda_{0})}\right)\left(\frac{2\pi}{3\beta+\lambda_{0}}\right)^{\frac{d}{2}}\leq\zeta_{0}\leq\exp\left(-f_{0}(0)+\frac{\|\nabla f_{0}(0)\|^{2}}{2(\beta+\lambda_{0})}\right)\left(\frac{2\pi}{\beta+\lambda_{0}}\right)^{\frac{d}{2}}.$
Since $f_{0}(0)=V_{0}(0)$, $\|\nabla f_{0}(0)\|=\|\nabla V_{0}(0)\|=\|\nabla V_{0}(0)-\nabla V_{0}(x^{\prime})\|\leq 3\beta\|x^{\prime}\|\lesssim\sqrt{\beta}$,
$1\leq\frac{\zeta_{0}}{\widehat{\zeta}_{0}}\leq\exp\left(\frac{\beta\|\nabla V_{0}(0)\|^{2}}{(\beta+\lambda_{0})(3\beta+\lambda_{0})}\right)\left(1+\frac{2\beta}{\beta+\lambda_{0}}\right)^{\frac{d}{2}}\leq\exp\left(\frac{\beta^{2}}{(\beta+\lambda_{0})(3\beta+\lambda_{0})}+\frac{d\beta}{\beta+\lambda_{0}}\right).$
So $\lambda_{0}\asymp\frac{d\beta}{\varepsilon}$ implies $\frac{\beta^{2}}{(\beta+\lambda_{0})(3\beta+\lambda_{0})}+\frac{d\beta}{\beta+\lambda_{0}}\lesssim\varepsilon$. ∎
Lemma 38.
For $k=K-1$, $\lambda_{k}\asymp\frac{\beta}{\sqrt{d}}$ implies Eq. 41.
Proof.
When $k=K-1$, $g_{k}=\exp\left(\frac{\lambda_{k}}{2}\|\cdot\|^{2}\right)$. We have
$\frac{\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}^{2}}{\left(\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}\right)^{2}}$ $=\operatorname{\mathbb{E}}_{\pi_{0}}\exp\left(\frac{\lambda_{k}}{2}\|\cdot\|^{2}\right)\operatorname{\mathbb{E}}_{\pi_{0}}\exp\left(-\frac{\lambda_{k}}{2}\|\cdot\|^{2}\right).$
Define
$h_{1}(\lambda):=\operatorname{\mathbb{E}}_{\pi_{0}}\exp\left(\lambda\|\cdot\|^{2}\right)\operatorname{\mathbb{E}}_{\pi_{0}}\exp\left(-\lambda\|\cdot\|^{2}\right),~{}\lambda\in\left[0,\frac{\beta}{4}\right].$
One can take derivative to obtain
$\frac{\mathrm{d}}{\mathrm{d}\lambda}\log h_{1}(\lambda)=\int_{-\lambda}^{\lambda}\operatorname{Var}_{\overline{\rho}_{s}}\|\cdot\|^{2}\mathrm{d}s,$
where $\overline{\rho}_{s}\propto\exp(-V_{0}+s\|\cdot\|^{2})$ is $\frac{\beta}{2}$-strongly-log-concave and thus satisfies $\frac{2}{\beta}$-LSI. Hence,
$\operatorname{Var}_{\overline{\rho}_{s}}\|\cdot\|^{2}\leq\frac{8}{\beta}\operatorname{\mathbb{E}}_{\overline{\rho}_{s}}\|\cdot\|^{2}.$
Let $x^{\prime}_{s}$ be the global minimizer of $V_{0}-s\|\cdot\|^{2}$. By Lem. 34, $\|x^{\prime}_{s}\|\leq R$. Leveraging Lem. 32, we have
$\operatorname{Var}_{\overline{\rho}_{s}}\|\cdot\|^{2}\lesssim\frac{1}{\beta}\left(\operatorname{\mathbb{E}}_{\overline{\rho}_{s}}\|\cdot-x^{\prime}_{s}\|^{2}+\|x^{\prime}_{s}\|^{2}\right)\leq\frac{1}{\beta}\left(\frac{2d}{\beta}+R^{2}\right)\lesssim\frac{d}{\beta^{2}}.$
So $\frac{\mathrm{d}}{\mathrm{d}\lambda}\log h_{1}(\lambda)\lesssim\frac{\lambda d}{\beta^{2}}$, and thus
$\frac{\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}^{2}}{\left(\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}\right)^{2}}=h_{1}\left(\frac{\lambda_{k}}{2}\right)=\exp\left(O\left(\frac{\lambda_{k}^{2}d}{\beta^{2}}\right)\right)=1+O\left(\frac{\lambda_{k}^{2}d}{\beta^{2}}\right)=1+O(1).$
Lemma 39.
For $k\in\left[\!\left[0,K-2\right]\!\right]$, Eq. 41 holds with $\gamma_{0}=\frac{1}{\sqrt{d}}$.
Proof.
One can write $g_{k}=\exp\left(\frac{\gamma_{0}\lambda_{k+1}}{2}\|\cdot\|^{2}\right)$. Simple calculation yields
$\frac{\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}^{2}}{\left(\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}\right)^{2}}$ $=\frac{\operatorname{\mathbb{E}}_{\pi_{0}}\exp\left(-\frac{(1+\gamma_{0})\lambda_{k+1}}{2}\|\cdot\|^{2}\right)\operatorname{\mathbb{E}}_{\pi_{0}}\exp\left(-\frac{(1-\gamma_{0})\lambda_{k+1}}{2}\|\cdot\|^{2}\right)}{{\operatorname{\mathbb{E}}_{\pi_{0}}\exp\left(-\frac{\lambda_{k+1}}{2}\|\cdot\|^{2}\right)}^{2}}.$
Define
$h_{2}(\gamma):=\operatorname{\mathbb{E}}_{\pi_{0}}\exp\left(-\frac{(1+\gamma)\lambda}{2}\|\cdot\|^{2}\right)\operatorname{\mathbb{E}}_{\pi_{0}}\exp\left(-\frac{(1-\gamma)\lambda}{2}\|\cdot\|^{2}\right),~{}\gamma\in\left[0,\frac{1}{2}\right].$
One can similarly show
$\frac{\mathrm{d}}{\mathrm{d}\gamma}\log h_{2}(\gamma)=\frac{\lambda^{2}}{4}\int_{1-\gamma}^{1+\gamma}\operatorname{Var}_{\widetilde{\rho}_{t}}\|\cdot\|^{2}\mathrm{d}t,$
where $\widetilde{\rho}_{t}\propto\exp\left(-V_{0}-\frac{t\lambda}{2}\|\cdot\|^{2}\right)$ is $(\beta+t\lambda)$-strongly-log-concave and thus satisfies $\frac{1}{\beta+t\lambda}$-LSI. Hence, $\operatorname{Var}_{\widetilde{\rho}_{t}}\|\cdot\|^{2}\leq\frac{8}{\beta+t\lambda}\operatorname{\mathbb{E}}_{\widetilde{\rho}_{t}}\|\cdot\|^{2}$.
Let $x^{\prime\prime}_{t}$ be the global minimizer of $V_{0}+\frac{t\lambda}{2}\|\cdot\|^{2}$. By Lem. 34, $\|x^{\prime\prime}_{t}\|\leq\frac{R}{1+\frac{t\lambda}{3\beta}}$. Therefore,
$\operatorname{Var}_{\widetilde{\rho}_{t}}\|\cdot\|^{2}\lesssim\frac{1}{\beta+t\lambda}\left(\operatorname{\mathbb{E}}_{\widetilde{\rho}_{t}}\|\cdot-x^{\prime\prime}_{t}\|^{2}+\|x^{\prime\prime}_{t}\|^{2}\right)\lesssim\frac{1}{\beta+t\lambda}\left(\frac{d}{\beta+t\lambda}+\frac{\beta^{2}R^{2}}{\left(\beta+t\lambda\right)^{2}}\right).$
As a result,
$\frac{\mathrm{d}}{\mathrm{d}\gamma}\log h_{2}(\gamma)$ $\lesssim\lambda^{2}\int_{1-\gamma}^{1+\gamma}\frac{1}{\beta+t\lambda}\left(\frac{d}{\beta+t\lambda}+\frac{\beta^{2}R^{2}}{\left(\beta+t\lambda\right)^{2}}\right)\mathrm{d}t$
$\leq\lambda^{2}\int_{1-\gamma}^{1+\gamma}\frac{1}{t\lambda}\left(\frac{d}{t\lambda}+\frac{\beta^{2}R^{2}}{t^{2}\lambda^{2}}\right)\mathrm{d}t$
$\lesssim\lambda^{2}\gamma\cdot\frac{1}{\lambda}\left(\frac{d}{\lambda}+\frac{\beta^{2}R^{2}}{\lambda^{2}}\right)=\gamma\left(d+\frac{\beta^{2}R^{2}}{\lambda}\right)$
$\implies\log\frac{h_{2}(\gamma_{0})}{h_{2}(0)}$ $\lesssim\gamma_{0}^{2}\left(d+\frac{\beta^{2}R^{2}}{\lambda}\right)=1+\frac{\beta^{2}R^{2}}{d\lambda}.$
Since $\lambda_{k+1}\geq\lambda_{K-1}\asymp\frac{\beta}{\sqrt[]{d}}$ and $R\lesssim\frac{1}{\sqrt[]{\beta}}$, $\frac{\beta^{2}R^{2}}{d\lambda_{k+1}}\lesssim 1$, so $\frac{\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}^{2}}{\left(\operatorname{\mathbb{E}}_{\rho_{k}}g_{k}\right)^{2}}\leq 1+O(1)$. ∎
Finally, one can compute the total complexity as follows. The choice $\lambda_{0}\asymp\frac{d\beta}{\varepsilon}$, $\lambda_{K-1}\asymp\frac{\beta}{\sqrt[]{d}}$, and $\lambda_{k}=\left(1+\frac{1}{\sqrt[]{d}}\right)\lambda_{k+1}$ implies $K=\widetilde{\Theta}(\sqrt{d})$, and thus $N\asymp\frac{K}{\varepsilon^{2}}=\widetilde{\Theta}\left(\frac{\sqrt{d}}{\varepsilon^{2}}\right)$. For each $k$, it is necessary to obtain $N$ i.i.d. approximate samples from $\rho_{k}$ that are $\delta\asymp\frac{1}{NK}=\widetilde{\Theta}\left(\frac{\varepsilon^{2}}{d}\right)$-close in TV distance. Using proximal sampler (Fan et al., 2023), the complexity for obtaining one sample is $\widetilde{O}(\sqrt{d})$ (note that the condition numbers of $f_{k}$’s are uniformly bounded by $3$), so the total oracle complexity is $NK\cdot\widetilde{O}(\sqrt{d})=\widetilde{O}\left(\frac{d^{\frac{3}{2}}}{\varepsilon^{2}}\right)$. $\square$
H.3 Path Integral Sampler and Controlled Monte Carlo Diffusion
In this section, we briefly discuss two learning-based samplers used for normalizing constant estimation and refer readers to the original papers for detailed derivations. The path integral sampler (PIS) shares structural similarities with the RDS framework discussed in Thm. 11, using the time-reversal of a universal noising process that transforms any distribution into a prior – such as the OU process in RDS that converges to the standard normal or the Brownian bridge in PIS that converges to the delta distribution at zero. In contrast, the controlled Monte Carlo diffusion (CMCD) extends the JE framework from Sec. 4, focusing on learning the compensatory drift term along an arbitrary interpolating curve $(\pi_{\theta})_{\theta\in[0,1]}$, as long as the density of each intermediate distribution $\pi_{\theta}$ is known up to a constant.
@@ H.3 Path Integral Sampler and Controlled Monte Carlo Diffusion
The PIS learns the drift term of a reference SDE that interpolates the delta distribution at 00 and the target distribution $\pi$, which is closely connected with the Brownian bridge and the Föllmer drift (Chewi, 2022).
@@
$u^{*}=\mathop{\mathrm{missing}}{argmin}_{u}\operatorname{KL}(\mathcal{Q}^{u}\|\mathcal{Q}^{*})\implies\mathcal{Q}^{u^{*}}=\mathcal{Q}^{*}.$
One can calculate the KL divergence between these path measures via Girsanov theorem (Lem. 1):
@@
which implies $Z=\operatorname{\mathbb{E}}_{\mathcal{Q}^{u}}{\mathrm{e}^{-W^{u}}}$, and $\operatorname{KL}(\mathcal{Q}^{u}\|\mathcal{Q}^{*})=\operatorname{\mathbb{E}}_{\mathcal{Q}^{u}}{W^{u}}+\log Z$. On the other hand, directly applying Lem. 1 gives
@@
In Zhang and Chen (2022, Theorem 3), the authors considered the effective sample size (ESS) defined by $\mathrm{ESS}^{-1}=\operatorname{\mathbb{E}}_{\mathcal{Q}^{u}}{\left(\frac{\mathrm{d}\mathcal{Q}^{*}}{\mathrm{d}\mathcal{Q}^{u}}\right)^{2}}$ as the convergence criterion, and stated that $\mathrm{ESS}\geq 1-\varepsilon$ as long as $\sup_{t\in[0,T]}\|u_{t}-u^{*}_{t}\|^{2}_{L^{\infty}}\leq\frac{\varepsilon}{T}$. However, this condition is generally hard to verify since the closed-form expression of $u^{*}$ is unknown, and the $L^{\infty}$ bound might be too strong. Using the criterion (Eq. 5) and the same methodology in proving the convergence of JE (Thm. 6), we can establish an improved result on the convergence guarantee of this estimator, relating the relative error to the training loss of $u$, which is defined as
@@
We borrow the notations from Sec. 4 due to its similarity with JE.
Given $(\widetilde{\pi}_{t})_{t\in[0,T]}$ and the ALD (Eq. 6), we know from the proof of Thm. 5 that to make $X_{t}\sim\widetilde{\pi}_{t}$ for all $t$, the compensatory drift term $(v_{t})_{t\in[0,T]}$ must generate $(\widetilde{\pi}_{t})_{t\in[0,T}$. Now, consider the task of learning such a vector field $(u_{t})_{t\in[0,T]}$ by matching the following forward and backward SDEs:
@@
where the loss is $\operatorname{KL}(\mathcal{P}^{\to}\|\mathcal{P}^{\leftarrow})$, discretized in training. Obviously, when trained to optimality, both $\mathcal{P}^{\to}$ and $\mathcal{P}^{\leftarrow}$ share the marginal distribution $\widetilde{\pi}_{t}$ at every time $t$. By Girsanov theorem (Lem. 15), one can prove the following identity for a.s. $X\sim\mathcal{P}^{\to}$: $\log\frac{\mathrm{d}\mathcal{P}^{\to}}{\mathrm{d}\mathcal{P}^{\leftarrow}}(X)=W(X)+C^{u}(X)-\Delta F$, where $\Delta F$ and $W(X)$ are defined as in Thm. 5, and
@@
The proof of this theorem follows the same reasoning as that of Prop. 40. For normalizing constant estimation,
$\operatorname{Pr}\left(\left|\frac{\widehat{Z}}{Z}-1\right|\right)=\mathcal{P}^{\to}\left(\left|\frac{\mathrm{d}\mathcal{P}^{\leftarrow}}{\mathrm{d}\mathcal{P}^{\to}}-1\right|\geq\varepsilon\right)\lesssim\frac{\operatorname{TV}(\mathcal{P}^{\to},\mathcal{P}^{\leftarrow})}{\varepsilon}\lesssim\frac{\sqrt{\operatorname{KL}(\mathcal{P}^{\to}\|\mathcal{P}^{\leftarrow})}}{\varepsilon}\lesssim 1.$

### v2

@@ H.1 Thermodynamic Integration
We first briefly review the thermodynamic integration (TI) algorithm. Its essence is to write the free-energy difference as an integral of the derivative of free energy. Consider the general curve of probability measures $(\pi_{\theta})_{\theta\in[0,1]}$ defined in ˜7. Then,
@@
One may choose time points $0=\theta_{0}<...<\theta_{M}=1$ and approximate ˜34 by a Riemann sum:
@@
where the expectation under each $\pi_{\theta_{\ell}}$ can be estimated by sampling from $\pi_{\theta_{\ell}}$. Nevertheless, there is a way of writing the exact equality instead of the approximation in ˜35: since
@@
We follow the notations used in Thm.˜3 to demonstrate the following claim: TI ˜36 is a special case of AIS with every transition kernel $F_{\ell}(x,\cdot)$ chosen as the perfect proposal $\pi_{\theta_{\ell}}$.
@@
and hence the AIS equality becomes the following identity, exactly the same as ˜34:
@@
(III) The distinction between equilibrium and non-equilibrium methods.
@@
H.2 Path Integral Sampler and Controlled Monte Carlo Diffusion
In this section, we briefly discuss two learning-based samplers used for normalizing constant estimation and refer readers to the original papers for detailed derivations. The path integral sampler (PIS) shares structural similarities with the RDS framework discussed in Thm.˜5, using the time-reversal of a universal noising process that transforms any distribution into a prior – such as the OU process in RDS that converges to the standard normal or the Brownian bridge in PIS that converges to the delta distribution at zero. In contrast, the controlled Monte Carlo diffusion (CMCD) extends the JE framework from Sec.˜3, focusing on learning the compensatory drift term along an arbitrary interpolating curve $(\pi_{\theta})_{\theta\in[0,1]}$, as long as the density of each intermediate distribution $\pi_{\theta}$ is known up to a constant.
@@ H.2 Path Integral Sampler and Controlled Monte Carlo Diffusion
The PIS learns the drift term of a reference SDE that interpolates the delta distribution at 0 and the target distribution $\pi$, which is closely connected with the Brownian bridge and the Föllmer drift (Chewi, 2022).
@@
$u^{*}=\mathop{\mathrm{argmin}}_{u}\operatorname{KL}(\mathcal{Q}^{u}\|\mathcal{Q}^{*})\implies\mathcal{Q}^{u^{*}}=\mathcal{Q}^{*}.$
One can calculate the KL divergence between these path measures via Girsanov theorem (Lem.˜2):
@@
which implies $Z=\operatorname{\mathbb{E}}_{\mathcal{Q}^{u}}{\mathrm{e}^{-W^{u}}}$, and $\operatorname{KL}(\mathcal{Q}^{u}\|\mathcal{Q}^{*})=\operatorname{\mathbb{E}}_{\mathcal{Q}^{u}}{W^{u}}+\log Z$. On the other hand, directly applying Lem.˜2 gives
@@
In Zhang and Chen (2022, Theorem 3), the authors considered the effective sample size (ESS) defined by $\mathrm{ESS}^{-1}=\operatorname{\mathbb{E}}_{\mathcal{Q}^{u}}{\left(\frac{\mathrm{d}\mathcal{Q}^{*}}{\mathrm{d}\mathcal{Q}^{u}}\right)^{2}}$ as the convergence criterion, and stated that $\mathrm{ESS}\geq 1-\varepsilon$ as long as $\sup_{t\in[0,T]}\|u_{t}-u^{*}_{t}\|^{2}_{L^{\infty}}\leq\frac{\varepsilon}{T}$. However, this condition is generally hard to verify since the closed-form expression of $u^{*}$ is unknown, and the $L^{\infty}$ bound might be too strong. Using the criterion (˜1) and the same methodology in proving the convergence of JE (Thm.˜2), we can establish an improved result on the convergence guarantee of this estimator, relating the relative error to the training loss of $u$, which is defined as
@@
We borrow the notations from Sec.˜3 due to its similarity with JE.
Given $(\widetilde{\pi}_{t})_{t\in[0,T]}$ and the ALD ˜2, we know from the proof of Thm.˜1 that to make $X_{t}\sim\widetilde{\pi}_{t}$ for all $t$, the compensatory drift term $(v_{t})_{t\in[0,T]}$ must generate $(\widetilde{\pi}_{t})_{t\in[0,T}$. Now, consider the task of learning such a vector field $(u_{t})_{t\in[0,T]}$ by matching the following forward and backward SDEs:
@@
where the loss is $\operatorname{KL}(\mathcal{P}^{\to}\|\mathcal{P}^{\leftarrow})$, discretized in training. Obviously, when trained to optimality, both $\mathcal{P}^{\to}$ and $\mathcal{P}^{\leftarrow}$ share the marginal distribution $\widetilde{\pi}_{t}$ at every time $t$. By Girsanov theorem (Lem.˜4), one can prove the following identity for a.s. $X\sim\mathcal{P}^{\to}$: $\log\frac{\mathrm{d}\mathcal{P}^{\to}}{\mathrm{d}\mathcal{P}^{\leftarrow}}(X)=W(X)+C^{u}(X)-\Delta F$, where $\Delta F$ and $W(X)$ are defined as in Thm.˜1, and
@@
The proof of this theorem follows the same reasoning as that of Prop.˜4. For normalizing constant estimation,
$\operatorname{Pr}\left(\left|\frac{\widehat{Z}}{Z}-1\right|\geq\varepsilon\right)=\mathcal{P}^{\to}\left(\left|\frac{\mathrm{d}\mathcal{P}^{\leftarrow}}{\mathrm{d}\mathcal{P}^{\to}}-1\right|\geq\varepsilon\right)\lesssim\frac{\operatorname{TV}(\mathcal{P}^{\to},\mathcal{P}^{\leftarrow})}{\varepsilon}\lesssim\frac{\sqrt{\operatorname{KL}(\mathcal{P}^{\to}\|\mathcal{P}^{\leftarrow})}}{\varepsilon}\lesssim 1.$